# Dynamic Conditional Optimal Transport through Simulation-Free Flows

### v1

@@ 5.1 Continuity Equation
We note that having $v_{t}$ be triangular is sufficient, but certainly not necessary, for the conditional continuity equation to almost surely hold. For instance, the vector field in $\mathbb{R}^{d}$ that rotates $\mathcal{N}(0,I)$ about the origin preserves all conditional distributions.
5.2 Absolutely Continuous Curves
We will show in Theorem 14 that every absolutely continuous curve $(\gamma_{t})\subseteq\mathbb{P}_{p}^{\mu}(Y\times U)$ is generated (in the sense of satisfying the continuity equation) by a triangular vector field such that ${\left\lVert v_{t}\right\rVert_{L^{p}(\gamma_{t},Y\times U)}\leq|\gamma^{\prime}|(t)}$ for almost every $t\in I$. This result is obtained as a generalization of the technique in Ambrosio et al. (2005, Theorem 8.3.1), which is an application of variational methods for PDEs.
We define the map $j_{q}:L^{q}(\gamma,Y\times U)\to L^{p}(\gamma,Y\times U)$ for $11/p+1/q=$ 1 via
@@ 5.2 Absolutely Continuous Curves
$\left\lVert j_{q}(w)\right\rVert^{p}_{L^{p}(\gamma,Y\times U)}=\left\lVert w\right\rVert_{L^{q}(\gamma,Y\times U}^{q}=\int_{Y\times U}\langle j_{q}(w),w\rangle\,\mathrm{d}\gamma(y,u).$ (69)
Theorem 14 (AC Curves in $\mathbb{P}_{p}^{\mu}(Y\times U)$).
Let $I\subset\mathbb{R}$ be an open interval, and suppose $\gamma_{t}:I\to\mathbb{P}_{p}^{\mu}(Y\times U)$ is an absolutely continuous in the $W_{p}^{\mu}$ metric with $|\gamma^{\prime}|(t)\in L^{1}(I)$. Then, there exists a Borel vector field $v_{t}(x)$ such that
  1. 1.
    $v_{t}$ is triangular, i.e. $v_{t}(y,u)=\iota(v_{t}^{U}(y,u))$ for some Borel $v_{t}^{U}:Y\times U\to U$
  2. 2.
    $v_{t}\in L^{p}(\gamma_{t},Y\times U)$ for a.e.-$t$
  3. 3.
    $\left\lVert v_{t}\right\rVert_{L^{p}(\gamma_{t},Y\times U)}\leq|\gamma^{\prime}|(t)$ for a.e.-$t$
  4. 4.
    $(v_{t},\gamma_{t})$ solve the continuity equation distributionally.
@@
Theorem 15 (Continuous Curves Generated by triangular Vector Fields).
@@
As a corollary of Theorem 14 and Theorem 15, we obtain a conditional version of the Benamou-Brenier theorem (Benamou and Brenier, 2000). We note that the proof of Theorem 16 largely follows the unconditional case (see e.g. Ambrosio et al. (2005, Chapter 8)), but we include it for the sake of completeness.
Theorem 16 (Conditional Benamou-Brenier).
Suppose $Y,U$ are separable Hilbert spaces, $1<p<\infty$, and $\mu\in\mathbb{P}(Y)$. Then, for any $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$, we have
[…]
@@
6 COT Flow Matching
Figure 1: Results for the Darcy flow inverse problem. We compare the posterior mean and variance of our method (COT-FM) to pCN (Cotter et al., 2013) and WaMGAN (Hosseini et al., 2023).
We have thus far seen that the COT problem (11) admits a dynamical formulation (99), where one may take the underlying vector fields to be triangular. We use these results to design a principled model for conditional generation based on flow matching (Lipman et al., 2022; Albergo et al., 2023b). More specifically, recent work (Tong et al., 2023; Pooladian et al., 2023a) has proposed to couple the source and target measures in flow matching via optimal transport. We extend these techniques to the conditional setting.
In this section we make several simplifying assumptions. First, we use the squared-distance cost $|(y_{0},u_{0})-(y_{1},u_{1})|^{2}$ throughout (i.e. $p=2$). Second, we assume that $Y$ and $U$ are finite dimensional Euclidean spaces, but we emphasize that our techniques may be applied to infinite dimensional spaces under additional regularity assumptions (Kerrigan et al., 2023, 2022; Lim et al., 2023; Baldassari et al., 2024). Third, we assume that our source and target measures admit densities with respect to the Lebesgue measure on $Y\times U$. This is sufficient (Prop. 4) for the existence of a unique COT plan which is induced by a triangular map. We will henceforth, in an abuse of notation, identify measures with their densities.
Flow Matching.
We assume that we have access to samples $z_{0}=(y_{0},u_{0})\sim\eta(y_{0},u_{0})\in\mathbb{P}_{p}^{\mu}(Y\times U)$ from a source measure, and samples $z_{1}=(y_{1},u_{1})\sim\nu(y_{1},u_{1})\in\mathbb{P}_{p}^{\mu}(Y\times U)$ from a target measure. Let $z=(z_{0},z_{1})\sim\rho(z_{0},z_{1})\in\Pi(\eta,\nu)$ be any coupling of the source and target measure. Following Tong et al. (2023), we specify a collection of measures and vector fields via
$\gamma_{t}(y,u\mid z)=\mathcal{N}\left(y,u\mid tz_{1}+(1-t)z_{0},C\right)$ (102)
$u_{t}(y,u\mid z)=z_{1}-z_{0}$ (103)
where $C$ is any trace-class covariance operator (Da Prato and Zabczyk, 2014), e.g. $C=\sigma^{2}I$.
As is standard in flow matching (Lipman et al., 2022; Tong et al., 2023; Kerrigan et al., 2023), we obtain from Equations (102) and (103) a marginal measure $\gamma_{t}(y,u)$ and vector field $u_{t}(y,u)$ satisfying the continuity equation via
$\gamma_{t}(y,u)=\int_{(Y\times U)^{2}}\gamma_{t}(y,u\mid z)\,\mathrm{d}\rho(z)$ (104)
$u_{t}(y,u)=\int_{(Y\times U)^{2}}u_{t}(y,u\mid z)\frac{\,\mathrm{d}\gamma_{t}(y,u\mid z)}{\,\mathrm{d}\gamma_{t}(y,u)}\,\mathrm{d}\rho(z).$ (105)
To model the intractable $u_{t}(y,u)$, we minimize the loss [^1] This is often called the conditional flow matching loss (Tong et al., 2023), which is not to be confused with the notion of conditioning that we focus on in this work.
$\mathcal{L}(\theta)=\mathbb{E}_{t,\rho(z),\gamma_{t}(y,u\mid z)}\left\lVert u^{\theta}(t,y,u)-u_{t}(y,u\mid z)\right\rVert^{2}$ (106)
which has the same gradient (with respect to $\theta$ as the MSE loss to the true vector field of interest $u_{t}(y,u)$ (Tong et al., 2023, Theorem 3.2).
COT Flow Matching.
Figure 2: Further visualizations of our results on the Darcy flow inverse problem. The first row shows the ground-truth posterior. Row two shows samples obtained from out COT-FM method, and row three shows the same for WaMGAN (Hosseini et al., 2023). The prior and conditioning information are shown in the final two rows.
In the preceding section, $\rho(z)$ may be an arbitrary coupling between $\eta$ and $\nu$. Motivated by Propositions 1 and 4, we will choose $\rho$ to be a conditional optimal coupling, so that $(y_{0},u_{0},y_{1},u_{1})\sim\rho$ has the form $\left(y,u,y,T_{U}(y,u)\right)$. By Theorem 16 and Theorem 11, under such a $\rho$ the vector fields in Equation (103) and Equation (105) will be triangular. Thus, we parametrize our model $u^{\theta}$ to also be triangular. If the covariance operator $C$ in Equation (102) is taken such that $\text{tr}(C)\to 0$, then we recover the optimal path of measures given by the conditional Benamou-Brenier Theorem via a pointwise application of (Tong et al., 2023, Proposition 3.4).
Given a collection of samples $\{z_{0}^{i},z_{1}^{i}\}_{i=1}^{n}$ drawn from $\eta$ and $\nu$, we approximate a conditional optimal coupling $\rho$ using standard numerical techniques (e.g., those available in the POT package (Flamary et al., 2021)) with the cost function
$c_{\epsilon}(y_{0},u_{0},y_{1},u_{1})=|y_{1}-y_{0}|^{2}+\epsilon|u_{1}-u_{0}|^{2}$ (107)
for some $0<\epsilon\ll 1$. Intuitively, such a cost penalizes mass transfer along the $Y$ dimension, which is precisely the constraint sought in the COT problem (11). Hosseini et al. (2023, Prop. 3.11) show that as $\epsilon\downarrow 0$, we recover the true optimal triangular map. When our datasets are not too large, we precompute this COT coupling prior to training our flow matching model. Otherwise, we compute this COT coupling for each minibatch drawn at training time.
After training, we obtain a learned triangular vector field $u^{\theta}_{t}(y,u)$. Given an arbitrary fixed $y\in Y$, we may approximately sample from the target $\nu(u\mid y)$ by sampling $u_{0}\sim\eta(u_{0}\mid y)$ and numerically solving the corresponding differential equation. For some tasks, like domain translation, note that we are given a fixed $u_{0}$ of interest rather than sampling $u_{0}$ from the source. In these cases, we are effectively sampling from $\nu(u\mid y,u_{0})$.
Source Measure.
Our framework is agnostic to the choice of source measure $\eta$, allowing for great flexibility in the modeling process. The main requirement is that the $Y$-marginal of $\eta$ must match the $Y$-marginal of the target $\nu$. In some scenarios, this is trivially satisfied. For instance, if one is interested in using a source distribution which is simply random noise, one may take $\eta(y_{0},u_{0})=\pi^{Y}_{\#}\nu(y_{0})\eta_{U}(u_{0})$ to be the product of two independent distributions where $\eta_{U}(u)$ is arbitrary, e.g. Gaussian noise. Sampling from such an $\eta$ can be achieved by sampling noise from $\eta_{U}(u)$ and an independent $y$ from the target measure. We explore more complex choices for $\eta$ in Section 7.
7 Experiments
The following section is devoted to illustrate our methodology on a variety of different problems related to conditional simulation. First, we explore a likelihood-free inference problem in function space on a Darcy flow inverse problem. The dataset used in this case mirrors those in Hosseini et al. (2023), as we aim to directly compare with their Conditional Wasserstein Monge GAN (WaMGAN) method. In addition, we demonstrate our method on two unpaired image-to-image translation tasks (Nazarpour and Chen, 2017; Choi et al., 2020)
7.1 Bayesian Inverse Problems
In this experiment, we consider simulation-based posterior inference on the parameters characterizing a stochastic forward model, whose likelihood evaluation is intractable.
Darcy Flow
In this experiment the forward model corresponding to the 2D Darcy flow, an ellipitic PDE on a smooth domain $\Omega\subset\mathbb{R}^{d}$ characterized by a permeability field $\exp(u)$, a pressure field $\rho$, and a source term $f$:
$-\text{div}\exp(u)\nabla\rho=f\quad\text{in}\quad\Omega,\qquad\rho=0\quad\text{on}\quad\partial\Omega.$ (108)
For simplicity we denote the deterministic part of the forward model as $\mathfrak{F}:U\rightarrow Y$, and the stochasticity arises from only having access to noisy measurements $y=\mathfrak{F}(u)+\epsilon\in Y,$ where $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$. Although the general setting of the experiment corresponds to the one in Hosseini et al. (2023), we consider the more general task of performing posterior inference for $u$ when both $u$ and $y$ are observed at arbitrary resolutions, effectively posing the problem in function space. Hence, we consider $Y\subseteq\mathcal{Y}$, $U\subseteq\mathcal{U}$, where $\mathcal{Y},\mathcal{U}$ are infinite-dimensional Hilbert spaces. We remark that, although the measurements $Y$ are observed on a grid their resolution does not need to be fixed, allowing for training on observations at different resolutions. It follows that $\epsilon$ follows a Gaussian measure on $Y$ with a white noise kernel. Notice that although the conditional probability measure $\nu^{u}:=\pi^{U}_{\#}\nu$ is Gaussian, evaluating it requires simulating from Equation (108).
Figure 3: Zero-shot super-resolution on the Darcy flow inverse problem. Our model (COT-FM) is trained using a Fourier Neural Operator (Li et al., 2020), allowing us to generate samples at higher resolutions than available at training time. We compare to WaMGAN (Hosseini et al., 2023) trained with the same architechture.
We assign to $u_{1}$ a prior corresponding to the Gaussian measure $\mathcal{N}(0,C_{1})$, where $C_{1}$ corresponds to a symmetric, non-negative and trace-class covariance operator (Da Prato and Zabczyk, 2014). In order to obtain a dataset $\mathcal{D}_{1}=\{y_{1,i},u_{1,i}\}_{i=1}^{n}$ of samples on the target domain, we simulate $y_{1}$ from Equation (108) parameterizing the forward model with samples $u_{1}\sim\mathcal{N}(0,C_{1})$. We denote the target measure as $\nu$. We construct the source measure $\eta$ as the product measure $\mathcal{N}(0,C_{0})\times\mu$, where $C_{0}$ is symmetric, non-negative, trace-class and $\mu=\pi^{Y}_{\#}\eta$ corresponds to the marginal data distribution. The choice of picking $C_{0}\neq C_{1}$, while complicating the problem, reflects a scenario where we want to perform inference with a weakly informative prior that does not reflect the true data generating process. Notice that sampling from a conditional coupling $\gamma\in\Pi_{Y}(\eta,\nu)$ would still recover samples from the true posterior. Here, we assume that $C_{0},C_{1}$ are sufficiently regular such that the path of measures in Equation (104) is well-defined.
@@ 7.1 Bayesian Inverse Problems
Figure 4: Samples from our model on a supervised image-to-image translation task, where translation is performed from Chinese MNIST to MNIST (a) and from MNIST to Chinese MNIST (b). This demonstrates that our method is able to use non-trivial source distributions and is able to successfully draw samples from the corresponding conditional measure.
The resulting amortized sampler, denoted for simplification by the mapping $(y,u_{0})\mapsto u_{1}=\tilde{T}_{U}(y,u_{0})$, will parameterize an approximate posterior measure. Notice that, in contrast to variational inference techniques, no distributional assumptions are made about the approximate posterior. In turn, integrals are obtained numerically by Monte Carlo sampling $K-$many samples from the prior, resulting in the approximation
@@
In Figure 1, we show an example of how our method compares to pCN and WaMGAN on recovering the posterior mean and variance on a given example. For a fair comparison and to ensure resolution-invariance, we used the same architecture for both neural models, with the discriminator having an additional layer projecting to a scalar value. However, the WaMGAN method used twice as many parameters as it requires two neural networks. Each statistic is computed on $2\times 10^{3}$ samples from each method. We showcase the ability of our method to perform zero-shot super-resolution in Figure 3. A random selection of samples from our method and from WaMGAN, as well as the prior sample and the different conditions used for generation, are displayed in Figure 2. Qualitatively, our samples appear to better match the true posterior compared to WaMGAN (Hosseini et al., 2023), while being cheaper to sample than pCN (Cotter et al., 2013).
7.2 Unpaired Image-to-Image Translation
Figure 5: Uncurated sample from the test set of the Animal Faces HQ dataset (Choi et al., 2020) from the dogs class, translated to cats. These results demonstrate our method on an unsupervised image-to-image task, where we utilize a pre-trained content extractor.
We now demonstrate our method on an image-to-image translation task. In particular, in this section our source measures are distributions over images. The objective of domain translation is to convert information from one domain to another by adjusting domain-specific attributes, denoted style, while preserving the domain-invariant ones, denoted content (Isola et al., 2017). Semantic attributes can be associated with human-assigned labels (supervised) or inferred by a learning algorithm that learns domain-invariant representations of the data (unsupervised). Pseudocode is available in Appendix A.
Historically, the problem of domain translation was tackled through methods that heavily relied on parallel datasets—pairs of corresponding samples from the source and target domains. However, obtaining such paired datasets is often impractical or impossible in many real-world applications, prompting the shift towards methods that do not require paired data (unpaired domain translation) (Pang et al., 2021). Domain translation is possible for a learning algorithm if it satisfies cycle consistency, positing that translation of a source sample $u_{s}$ into a target sample $\hat{u}_{t}$ and then back to the source domain should recover $\hat{u}_{s}=u_{s}$, and the same shall hold by starting from a target sample. Notice that this property is achieved naturally by our model, and hence does not require additional terms in the loss as is the case for GAN- or VAE-based methods (Zhu et al., 2017; Isola et al., 2017; Liu et al., 2017).
Here, the space $U$ corresponds to the space of images. For the space $Y$, we explore two choices. First, if supervised information (i.e., a class label) is available, we take $Y$ to be the space of labels. In this setting, if the class distribution in both the source and target distribution is balanced, then the $Y$-marginals of the source and target measures match. When such information is not available, we extract semantic information from our images using a pre-trained encoder, mapping images to a latent space $Y$. In this case, we assume the encoding distribution matches across the source and target distributions. Exploring techniques for balancing this distribution is an interesting avenue for future work.
Supervised.
We showcase in Figure 4 the ability of our method to address the task of class-conditional image translation, using the MNIST and the Chinese MNIST (Nazarpour and Chen, 2017) datasets. The second dataset was preprocessed in order to center the images around their content, and both are reshaped to a resolution of $32\times 32$. The size of the training set was small enough for us to compute the empirical COT plan on the entire dataset, and pairs were resampled at each epoch. We parameterize the neural network in COT-FM with a UNet (Ronneberger et al., 2015), and we compute the $Y-$distance in the COT plan by taking the $L_{2}$ distance between one-hot encoded vectors for the labels. The UNet has hidden dimensionality of 128, and we use three downsampling/upsampling blocks with multipliers $(1,2,2)$. The optimizer is Adam (Kingma and Ba, 2014) with a learning rate of $10^{-4}$, at a batch size of 128 for $3\times 10^{5}$ training steps.
Unsupervised.
Samples are mapped into a shared representation space by a pre-trained feature extractor $\mathfrak{F}:\mathbb{R}^{d}\rightarrow[0,1]^{e}$, obtaining content variables $y=\mathfrak{F}(u)$. These effectively act as pseudo-labels when building the empirical COT matrix. In our example, we perform translation from the dogs class to the cats class in the AnimalFacesHQ dataset (Choi et al., 2020) and we use CLIP as our feature extractor (Radford et al., 2021) as found in the official Hugging Face repository in its version clip-vit-large-patch14. The training set is comprised of 5067 images of cats and 4680 images of dogs, enabling the computation of the COT plan in advance. As with the previous experiment, samples are repaired by sampling from the COT plan at each epoch. Images are observed at a native resolution of $512\times 512$, but to ease the memory requirements we employ another pre-trained encoder $\mathfrak{E}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{z}$ for dimensionality reduction, effectively performing flow matching in latent space (Dao et al., 2023). The procedure is detailed in Algorithm 1. We take the encoder $\mathfrak{E}$ to be the variational autoencoder (Kingma and Welling, 2013) from Stable Diffusion v2.1, enabling training of the COT-FM network at a dimensionality of $4\times 64\times 64$. We use the Adam optimizer (Kingma and Ba, 2014) with a learning rate of $2\times 10^{-4}$, with a batch size of 32 for $4\times 10^{5}$ training steps. We show several uncurated conditional samples from our model in Figure 5.
8 Conclusion
We analyze conditional optimal transport from a geometric and dynamical point of view. Our analysis culminates in the characterization of absolutely continuous curves of measures in a conditional Wasserstein space, resulting in a conditional analog of the Benamou-Brenier Theorem. We use these result to build on the framework of triangular transport and flow matching to develop simulation-free methods for conditional generative models. Our methods are applicable across a wide class of problems, and we demonstrate our methodology on an infinite-dimensional Bayesian inverse problem as well as image translation tasks.
Acknowledgments
Appendix A Pseudocode
1:  Input: Data $\{u_{0,i},u_{1,i}\}_{i=1}^{B}\sim\gamma^{\star}_{COT}$, pretrained content extractor $\mathfrak{F}$, interpolant network $v_{\theta}$, pretrained encoder $\mathfrak{E}$ (optional).
2:  Initialize: $\mathcal{L}=0$
3:  $\{y_{0,i},y_{1,i}\}_{i=1}^{B}\leftarrow\{\mathfrak{F}(u_{0,i}),\mathfrak{F}(u_{1,i})\}_{i=1}^{B}$
4:  if  $\mathfrak{E}$ is given  then
5:     $\{u_{0,i},u_{1,i}\}_{i=1}^{B}\leftarrow\{\mathfrak{E}(u_{0,i}),\mathfrak{E}(u_{1,i})\}_{i=1}^{B}$
6:  end if
7:  for $i=1,\dots,B$ do
8:     Sample $t_{i}\sim\mathcal{U}([0,1])$
9:     Interpolate: $u_{t,i}=(1-t_{i})u_{0,i}+t_{i}u_{1,i}$
10:     Evaluate interpolant network:
$\hat{v}_{t_{i}}^{0}=v_{\theta}(u_{t,i},t_{i},y_{0,i}),\quad\hat{v}_{t_{i}}^{1}=v_{\theta}(u_{t,i},t_{i},y_{1,i})$
[…]
12:  end for
13:  $\theta_{\text{new}}\leftarrow$ OptimizerStep$(\nabla_{\theta}\mathcal{L})$
Algorithm 1 DCOT Flow for Unsupervised Image-to-Image Translation, Training Step

### v2

@@ E.1 Continuity Equation
E.2 Absolutely Continuous Curves
We now proceed to prove the main results of this section. First, we introduce some preliminary notions. We define the map $j_{q}:L^{q}(\gamma,Y\times U)\to L^{p}(\gamma,Y\times U)$ for $11/p+1/q=$ 1 via
@@ E.2 Absolutely Continuous Curves
$\left\lVert j_{q}(w)\right\rVert^{p}_{L^{p}(\gamma,Y\times U)}=\left\lVert w\right\rVert_{L^{q}(\gamma,Y\times U)}^{q}=\int_{Y\times U}\langle j_{q}(w),w\rangle\,\mathrm{d}\gamma(y,u).$ (71)
See also Ambrosio et al. (2005, Chapter 8).
See 3
@@
We now prove, in some sense, a converse of the previous theorem.
See 4
@@
Together, Theorem 3 and Theorem 4 give us a dynamical interpretation of the conditional Wasserstein distance. The following result is a conditional analogue of the well-known Benamou-Brenier Theorem (Benamou and Brenier, 2000). Here, we note that the following proof follows the standard proof closely – the main legwork in obtaining this conditional generalization is through the previous two theorems.
See 5
@@
Appendix F Experiment Details
In this section, we provide additional details regarding all of our experiments, as well as additional results not contained within the main paper. All models can be trained on a single GPU with less than 24 GB of memory, and our experiments were parallelized over 8 such GPUs on a local server. We first describe our setting for the 2D and Lotka-Volterra experiments, as these share a similar setup. Details for the Darcy flow inverse problem are described in the corresponding section.
Models.
For FM and COT-FM, our model architecture is an MLP with SeLU activations (Klambauer et al., 2017). Time conditioning is achieved by concatenating the time variable as an input to the network. The covariance operator $C$ chosen in the path of measures in Equation (8) is taken to be $C=\sigma^{2}I$ where $\sigma$ is a hyperparameter.
Our implementation of FM is adapted from the torchcfm package Tong et al. (2023), available under the MIT License. For PCP-Map and COT-Flow, we adapt the open-source implementations from Wang et al. (2023), available under the MIT License.
Training and Model Selection.
Hyperparameter tuning of the PCP-Map and COT-Flow models was performed directly using the code of Wang et al. (2023), essentially implementing grid-search with an early stopping procedure. We refer to the paper and codebase of Wang et al. (2023) for further details. For COT-FM and FM, we perform a random grid search over 100 hyperparameter settings using the grid described in Table 4. For all model types, we select the best model used to generate the results in the paper as the training checkpoint that resulted in the lowest $W_{2}$ error to the joint target distribution on a held-out validation set. For training, we use the Adam optimizer where we only tune the learning rate, leaving all other settings as their defaults in pytorch.
Table 4: Hyperparameter grid used for random search of the FM and COT-FM models on the 2D and Lotka-Volterra datasets.
Hyperparameter Description Values
$\epsilon$ COT coupling strength [1e-6, 1e-4, 1e-2, 1e-1]
$\sigma$ Variance for $C=\sigma^{2}I$ in (8) [1e-3, 1e-2, 1e-1, 5e-1]
Batch Size Training batch size [256, 512, 1024]
Width Layer width in MLP [256, 512, 1024, 2048]
LR Learning rate [1e-4, 3e-4, 7e-4, 1e-3]
Layers Number of MLP layers [4, 6, 8]
F.1 2D Synthetic Data
Data Generation.
This experiment consists of four 2D synthetic datasets, where $Y=U=\mathbb{R}$. The datasets moons, circles, swissroll are available through scikit-learn (Pedregosa et al., 2011). The moons dataset is generated with noise=$0.05$ followed by standard scaling with a mean of $m=(0.5,0.25)$ and standard deviation of $\sigma=(0.75,0.25)$. The circles dataset is generated with factor=0.5 and noise=0.05. The swissroll dataset is generated with noise=0.75, followed by projection to the first two coordinates and re-scaling by a factor of $12$. All other unstated parameters are left as their default values. We use the code available from Hosseini et al. (2023) to generate the checkerboard dataset. For all datasets, we generate a training set (i.e., samples from the target distribution) of $20,000$ samples and $1,000$ held-out validation samples for model selection. Means and standard deviations in Table 1 are reported across five independent testing sets of 5,000 samples for the best representative of each model type.
In COT-FM, to generate samples from the source distribution, we sample an additional $20,000$ points from the target distribution and keep only the $Y$ coordinates. This ensures that the source and target have equal $Y$ marginals. During training, standard Gaussian noise $\mathcal{N}(0,1)$ is sampled for the $U$ coordinate of these source points at each minibatch.
We use minibatch COT couplings (Tong et al., 2023) in this experiment as computing the full COT plan was prohibitively expensive in terms of memory usage. However, we note that we use large batch sizes, meaning that the COT plan we find in this way should not be too far from optimal. All couplings are computed using the POT Python package (Flamary et al., 2021).
Figure 5: Samples from the ground-truth joint target distribution and the various models for the 2D datasets. Samples from COT-FM more closely match the ground-truth distribution than the baselines. A common failure mode for the baselines is to generate samples from regions with zero support under the true data distributions. Table 1 contains a quantitative evaluation.
Figure 6: Conditional KDEs shown for each of the methods on the 2D datasets. The conditioning variable $y$ is fixed at the horizontal dashed line shown in Figure 5. In all plots, the orange solid line indicates the CKDE of the ground-truth joint samples. In each column, the dashed blue line indicates the CKDE of samples generated from the respective method.
F.2 Lotka-Volterra Dynamical System
Data Generation.
We adopt the settings of Alfonso et al. (2023) for this experiment. As described in the main paper, we assume $p(0)=(30,1)$ and that $\log(u)\sim\mathcal{N}(m,0.5I)$ with $m=(-0.125,-3,-0.125,-3)$. Given parameters $u\in\mathbb{R}^{4}_{\geq 0}$, we simulate Equation (11) for $t\in\{0,2,\dots,20\}$ to obtain a solution $z(u)\in\mathbb{R}^{22}_{\geq 0}$. An observation $y\in\mathbb{R}^{22}_{\geq 0}$ is obtained by the addition of log-normal noise, i.e. $\log(y)\sim\mathcal{N}(\log(z(u),0.1I)$. We thus may simulate many $(y,u)$ pairs from the target measure for training.
We generate a training set of $10,000$ $(y,u)$ pairs using the procedure described above and a held-out validation set of $10,000$ $(y,u)$ pairs for model selection. Means and standard deviations in Table 2 are reported across five independent testing sets of 5,000 samples for the best representative of each model type. Figure 2 and Figures 9, 8, 7, 10 show $10,000$ samples from each model, as well as $10,000$ samples from the differential evolution Metropolis MCMC sampler (Braak, 2006) after a burn-in of $50,000$ samples. This is implemented through the PyMC Python package (Abril-Pla et al., 2023).
For COT-FM we use the full COT couplings, i.e. without minibatches. This is available to use due to the smaller size of the training set used in this experiment. The COT couplings are computed in the same way as the previous section, and as described in Section 7.
Figure 7: KDE plots of the samples on the Lotka-Volterra system, using the settings described in Section 7. Plots include one-dimensional KDEs on the diagonal, as well as all two-dimensional pairs. In all plots, samples from MCMC are drawn in orange, and samples from our method (COT-FM) are indicated in blue. The true unknown parameters are indicated by the red vertical line in the diagonal plots, or the black x in the off-diagonal plots.
Figure 8: KDE plots of the samples on the Lotka-Volterra system, using the settings described in Section 7. Plots include one-dimensional KDEs on the diagonal, as well as all two-dimensional pairs. In all plots, samples from MCMC are drawn in orange, and samples from PCP-Map are indicated in blue. The true unknown parameters are indicated by the red vertical line in the diagonal plots, or the black x in the off-diagonal plots.
Figure 9: KDE plots of the samples on the Lotka-Volterra system, using the settings described in Section 7. Plots include one-dimensional KDEs on the diagonal, as well as all two-dimensional pairs. In all plots, samples from MCMC are drawn in orange, and samples from COT-Flow are indicated in blue. The true unknown parameters are indicated by the red vertical line in the diagonal plots, or the black x in the off-diagonal plots.
Figure 10: KDE plots of the samples on the Lotka-Volterra system, using the settings described in Section 7. Plots include one-dimensional KDEs on the diagonal, as well as all two-dimensional pairs. In all plots, samples from MCMC are drawn in orange, and samples from flow matching (FM) are indicated in blue. The true unknown parameters are indicated by the red vertical line in the diagonal plots, or the black x in the off-diagonal plots.
F.3 Inverse Darcy Flow
Dataset.
The training and test datasets are generated following the same procedure as Hosseini et al. (2023): pressure fields $u$ are sampled from a Gaussian process with Matérn kernel having $\nu=3/2$ and lengthscale $\ell=1/2$, on a regular $40\times 40$ grid. The parameters are then exponentiated and used to simulate the permeability fields $p$ from the forward model $\mathfrak{F}$ solving the Darcy flow PDE, using FEniCS (Alnæs et al., 2015). Stochasticity arises from adding Gaussian noise to the permeability fields, obtaining $y=\mathfrak{F}(u)+\epsilon,\,\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$. For our experiments we observe $y$ on a $100\times 100$ grid, and we use $\sigma=2.5\times 10^{-2}$. We note that this level of noise is quite considerable, as it accounts for roughly $60\%$ of the variability in the $y$. Figure 11 showcases a data point for reference. Our source and target training sets contain $1\times 10^{4}$ samples each, and our test set comprises $5\times 10^{3}$ samples. We remark that although $y$ and $u$ are observed on a grid their resolution does not need to be fixed, allowing for training at different resolutions.
Figure 11: Example of one random data point from the Darcy flow dataset.
Figure 12: Additional results on the Darcy flow inverse problem. A true permeability $u$ is shown, as well as the pressure field $\rho$ and its observed, noisy version $y$. We compare an ensemble average of posterior samples from the various methods against MCMC (pCN) (Cotter et al., 2013). COT-FM achieves the lowest MSE to pCN. We note here that WaMGAN has clear visual artifacts despite achieving reasonable MSE and CRPS scores.
Models.
@@ F.3 Inverse Darcy Flow
  • WaMGAN (Hosseini et al., 2023): using an adaptation to the FNO architecture of the original code [^3] https://github.com/TADSGroup/ConditionalOT2023, we perform a grid search as detailed in Table 5. We found the training procedure to be rather unstable, and for this reason we checkpoint the model every 100 epochs and report the results for the best performing model at its best checkpoint. We found this to be a model with learning rate $1\times 10^{-4}$, 2 full critic iterations, and monotone penalty of $1\times 10^{-3}$. The gradient penalty parameter did not seem to significantly affect performance on the test set, and was set to 5.
  • FFM (Kerrigan et al., 2024): the learning rate is fixed to $5\times 10^{-4}$, and the covariance operator $C$ is set to match that of the prior, but rescaled by a factor of $\sigma=1\times 10^{-3}$. We use the code from the original repository [^4] https://github.com/GavinKerrigan/functional_flow_matching.
  • COT-FFM: we set $\epsilon=1\times 10^{-5}$ in the cost function used to build the COT plan. The learning rate and $C$ are chosen to be the same as FFM. In order to build COT couplings, we take the source measure to be the product measure $\pi^{Y}_{\#}\eta\times\mathcal{N}(0,C)$. Approximate couplings are obtained on minibatches of size $256$.
It should be noted that in any scenario where the source and the target $U-$marginals are identical, using the OT coupling would yield the identity mapping as the optimal vector field minimizing (10). Hence, the OT-CFM model (Tong et al., 2023) is inapplicable here.
Sampling.
The resulting amortized sampler, denoted for simplification by the mapping $(y,u_{0})\mapsto u_{1}=\tilde{T}_{U}(y,u_{0})$, will parameterize an approximate posterior measure. Notice that, in contrast to classical variational inference techniques, no distributional assumptions are made about the approximate posterior. In turn, integrals are obtained numerically by Monte Carlo sampling $K$ samples from the prior, resulting in the approximation
@@
Table 5: Hyperparameter search space for WaMGAN
Parameter Search Space
Learning rate $1\times 10^{-3}$, $5\times 10^{-4}$, $1\times 10^{-4}$}
Full critic iter. {2, 5, 10}
Monotone penalty $1\times 10^{-3}$, $5\times 10^{-2}$, $1\times 10^{-1}$}
Gradient penalty {1, 5, 10}